OpenStudy (anonymous):

can anyone say how to solve a differential equation using method of undetermined coefficients??

OpenStudy (anonymous):

give me the equation

OpenStudy (anonymous):

ok( D^2-5D+6)y=e^3x+sinx

OpenStudy (anonymous):

just tell me the assumed particular integral.....

OpenStudy (anonymous):

bad diff equation, what is D?

OpenStudy (anonymous):

D is the notation for operator d/dx....

OpenStudy (anonymous):

\[y^{\prime\prime}-5y^\prime+6y=\operatorname{e}^{3x}+\sin x\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I think one of the best ways to solve this equation is by using Laplace transformation

OpenStudy (anonymous):

in the question it is specifically mentioned to use method of undetermined coefficients

OpenStudy (anonymous):

The particular solution will be the sum of particular solutions. Try Ae^(3x) + Bsin(x)

OpenStudy (anonymous):

solve for A and B

OpenStudy (anonymous):

no it will not work as e^3x is a term in complementary function

OpenStudy (anonymous):

particular solution \[y=Ae^{\lambda1 x}+Be^{\lambda2 x}\], lambda are complex numbers

OpenStudy (anonymous):

yeah...i just did it and it's a mess...too much to write out...buggin' out and going to bed...good luck.

OpenStudy (anonymous):

I thought the others were going to help you. Since e^(3x) is part of the homogeneous solution, try xAe^(3x) as one part. Since you have a trig. function too, add to this Msin(x) + Ncos(x). So try,\[y_p=Axe^{3x}+M \sin x +N \sin x\]

OpenStudy (anonymous):

I think I ended up with A=1, M=N=(1/10). Try it.

OpenStudy (anonymous):

\[C _{1} e^{2x} + C_{2}e^{3x} +0.1(\cos[x] + \sin[x])+xe^{3x}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanx i ended up with the same solution guess the answer in the book was wrong